Please help me with this problem! Find the number of the sides of a regular polygon if each exterior angle measure is 18?? I'm stumped. : Please help me with this problem! Find the number of the sides of a regular polygon if each exterior angle measure is 18?? I'm stumped.
Answer by Alan3354(1944)

Please help me with this problem! Find the number of the sides of a regular polygon if each exterior angle measure is 18?? I'm stumped.
-------------------------
The sum of the INTERIOR angle of a polygon with n sides is 180*(n-2). Each angle is (180*(n-2))/n.
The exterior angles are 180-interior angle, so this is a polygon with interior angles of 162 degrees.
-------------
So, 180*(n-2)/n = 162
180*(n-2) = 162n
180n-360 = 162n
18n = 360
n = 20
I don't remember the name of that one, maybe duodecagon.